# Continuum Attention for Neural Operators

### v3

@@ 1 Introduction
Attention was introduced in the context of neural machine translation in Bahdanau et al. (2014) to effectively model correlations in sequential data. Many attention mechanisms used in conjunction with recurrent networks were subsequently developed, for example long short-term memory (Hochreiter and Schmidhuber, 1997) and gated recurrent units (Chung et al., 2014), which were then employed for a wide range of applications, see for example the review Chung et al. (2014). The seminal work Vaswani et al. (2017) introduced the now ubiquitous transformer architecture, which relies solely on the attention mechanism and not on the recurrence of the network in order to effectively model nonlocal, long-range correlations. In its full generality, the attention function described in Vaswani et al. (2017) defines an operation between sequences $\{u_{i}\}_{i=1}^{N}\subset\mathbbm{R}^{d_{u}},\{v_{i}\}_{i=1}^{M}\subset\mathbbm{R}^{d_{v}}$ of lengths $N,M\in\mathbbm{N}$ respectively, represented as tensors $u\in\mathbbm{R}^{N\times d_{u}},v\in\mathbbm{R}^{M\times d_{v}}$, where learnable weights $Q\in\mathbbm{R}^{d_{K}\times d_{u}},K\in\mathbbm{R}^{d_{K}\times d_{v}},V\in\mathbbm{R}^{d_{V}\times d_{v}}$ parametrize
@@ 1.1 Literature Review
The attention mechanism, as introduced in Vaswani et al. (2017) and defined in (1), has shown widespread success for a variety of tasks. In particular, given the suitability of attention and transformers for applications involving sequential data, there has been a large body of work extending the methodology from natural language processing, the setting in which it was originally introduced, to the more general context of time series. Namely, as surveyed in Wen et al. (2023), transformers and variants thereof have been applied to time series forecasting, classification and anomaly detection tasks. This generalization has involved modifications and innovations relating to the positional encoding and the design of the attention function itself. We refer the reader to Wen et al. (2023) for a detailed review of the large body of work and architectural variants of transformers that have been developed for time series applications.
@@
Attention in Vision The quadratic scaling in the input of the attention mechanism has made transformers a computationally prohibitive architecture for applications involving long sequences. Nonetheless, through techniques such as patching, the attention mechanism and transformers have also enjoyed success in computer vision, where high-resolution image data presents an increased computational cost. In fact, in vision transformers (ViT), introduced in Dosovitskiy et al. (2021), the input image is subdivided into patches; attention is then applied across the shorter sequence of patches, thus decreasing the overall cost.
@@
Neural Operators Given the success of the methodology in computer vision, there has been growing interest in extending the transformer methodology to the operator learning context. This operator setting concerns learning mappings between infinite dimensional spaces of functions. In this case, the domains on which the functions are defined (spatial and temporal) may inherently be a continuum. Neural operators (Kovachki et al., 2023), generalizations of neural networks that learn operators mapping between infinite-dimensional spaces, have been developed for this task. As they parametrize operators acting on spaces of functions, the learned parameters of neural operators may be applied to any discretization of the input function; this property of neural operators is referred to as “discretization invariance”. Discretization invariance is desirable both at a theoretical and practical level. Indeed, for a neural network to approximate an operator acting on functions defined on a continuum, its parameters cannot depend on the discretizations of the input functions that are used for learning the parameters themselves. On the other hand, discretization invariance allows to scalably train models on coarser discretizations of functions and to deploy or finetune trained models at different resolutions for downstream tasks. A recent example of the success of this approach is the GenCast medium-range weather forecasting model (Price et al., 2025), which because of discretization invariance properties is pretrained on low resolutions and finetuned at the desired predictive resolution.
@@
Neural Operators. An integral kernel interpretation of the attention mechanism appearing for example in Tsai et al. (2019); Martins et al. (2020); Moreno et al. (2022); Wright and Gonzalez (2021); Cao (2021); Kovachki et al. (2023) inspired a range of attention-based architectures for operator learning. For example, the work of Cao (2021) leverages insight from the integral kernel formulation to design novel attention mechanisms that do not employ the softmax function. The author designs attention mechanisms using integral kernels which may be approximated by a Fourier-type projection, leading to quadratic complexity, and by a Galerkin projection, leading to linear complexity. While the resulting attention maps are discretization invariant, the full architecture proposed uses convolutional neural networks to downsample the input to a feature map of sufficiently small resolution for the method to be scalable. Hence the full architecture is not a neural operator. Furthermore, the lack of the softmax nonlinearity leads to a mechanism lacking the probabilistic description outlined in this paper, thus differing from the original definition of attention in Vaswani et al. (2017). In Li et al. (2024) the authors take the Fourier attention approach as in Cao (2021) to design a neural operator for large eddy simulation, while in Luo et al. (2024) the authors take the Galerkin attention approach from Cao (2021) to design a novel neural operator architecture. In Li et al. (2023) the authors present “OFormer” a neural operator with an encoder-decoder structure that employs cross-attention to obtain the latent embeddings and then stacked self-attention blocks; the model uses the attention from Cao (2021) and random Fourier features for query coordinate encodings while using a recurrent multi-layer perceptron for unrolling in time-dependent problems as opposed to masked attention in Vaswani et al. (2017). In the recent work Rahman et al. (2024), the authors propose a neural operator architecture that employs an attention mechanism that acts across channels of the function in the latent embedding, and do so in the context of pre-training a foundation model for PDE solution operators. We demonstrate how such an architecture can be derived from our formulation in Remark 16. As in our paper, various attention-related concepts are defined in a function-space setting, but our formulations are more general; this issue is discussed in detail at the relevant point in the paper. We note that the architectures introduced in this paper rely solely on the attention mechanism defined in the continuum, hence they are neural operators; furthermore, because of the simple design, with a slight modification to our architectures we may prove the first universal approximation theorem for transformer operators.
@@
Non discretization invariant architectures. In Guibas et al. (2022) the authors propose a modification of the Fourier neural operator that involves mixing of tokens in Fourier space. This is achieved by first applying the FFT to the sequence of tokens and then applying a MLP to each token, before inverting back to the original space. This approach differs fundamentally from the original definition of attention in Vaswani et al. (2017) and lacks the probabilistic description outlined in this paper. Furthermore, because of learnable positional embeddings and patching via convolution, the AFNO is not discretization invariant. Various other approaches have been developed. The work Hao et al. (2023) introduces “GNOT”, an architecture employing a normalized attention layer that handles multiple inputs. This design, along with a learned gating mechanism for input coordinates, allows the model to be applied to irregular meshes and multi-scale problems; however, this architecture employs an attention mechanism that does not involve rescaling by the grid, hence not directly allowing for discretization invariance. In Ovadia et al. (2023) the authors propose to use an architecture composed of a ViT in the latent space of a U-Net (Ronneberger et al., 2015). The proposed model takes as input the input field evaluated on a grid with associated coordinate values and outputs an approximation of the solution to PDE inverse problems. While it is shown that the architecture achieves state-of-the-art accuracy, the parameters in the scheme are not invariant to the input function resolution. In Wang et al. (2025) the authors design “Continuous ViT”, an architecture consisting of a ViT encoder and a cross-attention decoder. The method handles input functions defined on space and time and the method is shown to present evidence of discretization invariance, with test discretizations close to training discretizations. However, while the architecture is discretization invariant in time, it is not by design in space. This is because the tokenization procedure employed by standard ViTs is not discretization invariant; we elaborate on this point in Remark 15.
@@
Transformers for Continuum Data The use of attention-based transformer models in the context of dynamical systems and data assimilation problems, where spatial and temporal domains of the systems are continuous, is at its infancy. For a comprehensive review on the application of machine learning to data assimilation we refer the reader to Cheng et al. (2023). In this context, the efficacy of transformers has been explored for forecasting in numerical weather prediction. In Chattopadhyay et al. (2022), a transformer module is integrated in the latent space of a U-Net (Ronneberger et al., 2015) architecture in order to leverage the permutation equivariance property of the transformer to improve physical consistency and forecast accuracy. The work in Bi et al. (2023) achieves state-of-the-art performance among end-to-end learning methods for numerical weather prediction by applying a ViT-like architecture and employing the shifted window methodology from Liu et al. (2021). These approaches, however, are not discretization invariant by design.
@@
Mathematics of Transformers Mathematical foundations for the methodology are only now starting to emerge, however most analysis has been preformed in the finite-dimensional setting. In the recent work Geshkovski et al. (2023), the authors study the attention dynamics by viewing tokens as particles in an interacting particle system. The formulation of the continuous-time limit of the dynamics, where the limit is taken in “layer time” allows to investigate the emergence of clusters among tokens. In Yun et al. (2020), the authors prove that transformers are universal approximators of continuous permutation equivariant sequence-to-sequence functions with compact support. More broadly, developing a firm theoretical framework for the subject is necessary to understand the empirical performance of the methodology and to inform the design of novel attention-based architectures.
@@ 2 Continuum Attention
In this section we formulate a continuum analogue of the attention mechanism described in Vaswani et al. (2017). We note that in the transformer architecture of Vaswani et al. (2017), a distinction is made between the self-attention and cross-attention functions. The former, a particular instance of (1) where $u=v$, can be used to produce a representation of the input that captures intra-sequence correlations. On the other hand, cross-attention is used to output a representation of the input $u$ that models cross-correlations between both inputs $u$ and $v$. Indeed, in Vaswani et al. (2017) cross-attention is employed in the decoder component of the transformer architecture so that the input to the decoder attends to the output of the encoder. For natural language applications in particular, the form of the cross-attention function has allowed for sequences of arbitrary length $N$ to be outputted from a transformer architecture while attending to sequences of different lengths, e.g. as outputs of length $M$ from a self-attention encoder block.
For pedagogical purposes we subdivide the presentation between the self-attention operator and the cross-attention operator, which are treated in Subsection 2.1 and Subsection 2.2 respectively. In Subsection 2.1.1 we define self-attention for discrete sequences indexed on finite bounded sets. Here, we outline the interpretation of self-attention as an expectation under a family of carefully defined discrete probability measures, each acting on the space of sequence indices. We show that under this formulation we recover the standard definition of self-attention from Vaswani et al. (2017). We proceed in Subsection 2.1.2 by formulating self-attention as an operator mapping between spaces of functions defined on bounded uncountable sets. This interpretation is natural, as functions may be thought of as “sequences” indexed over an uncountable domain. As in the discrete case, we formulate the self-attention operator as an expectation under a carefully defined probability measure acting on the domain of the input. We obtain an approximation result given by Theorem 6 that shows that for continuous functions the self-attention mechanism as implemented in practice may be viewed as a Monte Carlo approximation of the self-attention operator described. We mimic these constructions in Subsections 2.2.1 and 2.2.2 for the discrete and continuous formulations of cross-attention, respectively, and obtain an analogous approximation result given by Theorem 12. As in (1), throughout this section $Q\in\mathbbm{R}^{d_{K}\times d_{u}},K\in\mathbbm{R}^{d_{K}\times d_{v}},V\in\mathbbm{R}^{d_{V}\times d_{v}}$ are learnable weights that parametrize the attention operators. When we consider self-attention we will have $d_{v}=d_{u}.$
@@ 3 Continuum Patched Attention
[…]
@@ 3.1 Patched Self-Attention
An image may be viewed as a mapping $u:D\to\mathbbm{R}^{d_{u}}$, where $D\subset\mathbbm{Z}^{2}$, and where $d_{u}=1$ for grey-scale and $d_{u}=3$ for RGB-valued images. In the context of vision transformers (Dosovitskiy et al., 2021), the input to the attention mechanism is a sequence of $P$ “flattened” patches of the input image; we next outline the details to this procedure.
Letting $N$ be the total number of discretization points in $D$ and $P$ the number of patches, in the discrete setting each function patch $\widetilde{u}(p):D^{\prime}\to\mathbbm{R}^{d_{u}}$ may be represented as an $\mathbbm{R}^{N/P\times d_{u}}$ matrix. The flattening procedure of each patch in ViT involves reshaping this matrix into an $\mathbbm{R}^{N/P\cdot d_{u}}$ vector. The full image is thus represented as a sequence of $P$ vectors of dimension $N/P\cdot d_{u}$. Each patch vector is then linearly lifted to an embedding space of dimension $d_{\textrm{model}}$. The embedded image, to which attention is applied, may hence be viewed as a matrix $\widetilde{u}\in\mathbbm{R}^{P\times d_{\textrm{model}}}$. Therefore, ViT may be cast in the setting of the discrete interpretation of the self-attention operator outlined in Subsection 2.1.1, where the domain on which attention is applied is $D^{P}$. A key observation to the ensuing discussion is that the application of a linear transformation to a “flattened” patch in ViT breaks the mesh-invariance of the architecture; however, this operation may be viewed as a nonlocal transformation on each patch. We leverage this insight in Section 4 to design discretization invariant transformer neural operators that employ patching. We further note that in practice, patching in ViT is achieved via strided convolutions, with stride equivalent to the kernel size; however, because the kernel size determines the number of parameters, such a procedure also breaks discretization invariance as it leads to architectures with parameters dependent on resolution of the input.
@@
The work of Rahman et al. (2024) employs an attention mechanism that acts across the channels of the function $u\in\mathcal{U}\bigl{(}D;\mathbbm{R}^{d_{\textrm{model}}}\bigr{)}$ in the latent space of a neural operator architecture. Indeed the “codomain” attention mechanism defined may be viewed as a particular variant of Definition 14, where $D^{\prime}\cong D$, where $\mathsf{A}_{\mathsf{P}}(\widetilde{u})(j)$ is defined for $j\in\llbracket 1,d_{\textrm{model}}\rrbracket$ and where $\mathsf{Q},\mathsf{K},\mathsf{V}$ are implemented as integral operators. In particular the indexing is by channels and not by patch. For problems in computer vision, similar ideas have been explored with discrete versions of the $\mathsf{Q},\mathsf{K},\mathsf{V}$ operators (Chen et al., 2017).

### v4

@@ 1 Introduction
Attention was introduced in the context of neural machine translation in bahdanau2014neural to effectively model correlations in sequential data. Many attention mechanisms used in conjunction with recurrent networks were subsequently developed, for example long short-term memory (hochreiter1997long) and gated recurrent units (chung2014empirical), which were then employed for a wide range of applications, see for example the review chung2014empirical. The seminal work vaswani2017attention introduced the now ubiquitous transformer architecture, which relies solely on the attention mechanism and not on the recurrence of the network in order to effectively model nonlocal, long-range correlations. In its full generality, the attention function described in vaswani2017attention defines an operation between sequences $\{u_{i}\}_{i=1}^{N}\subset\mathbbm{R}^{d_{u}},\{v_{i}\}_{i=1}^{M}\subset\mathbbm{R}^{d_{v}}$ of lengths $N,M\in\mathbbm{N}$ respectively, represented as tensors $u\in\mathbbm{R}^{N\times d_{u}},v\in\mathbbm{R}^{M\times d_{v}}$, where learnable weights $Q\in\mathbbm{R}^{d_{K}\times d_{u}},K\in\mathbbm{R}^{d_{K}\times d_{v}},V\in\mathbbm{R}^{d_{V}\times d_{v}}$ parametrize
@@ 1.1 Literature Review
The attention mechanism, as introduced in vaswani2017attention and defined in (1), has shown widespread success for a variety of tasks. In particular, given the suitability of attention and transformers for applications involving sequential data, there has been a large body of work extending the methodology from natural language processing, the setting in which it was originally introduced, to the more general context of time series. Namely, as surveyed in wen2023transformers, transformers and variants thereof have been applied to time series forecasting, classification and anomaly detection tasks. This generalization has involved modifications and innovations relating to the positional encoding and the design of the attention function itself. We refer the reader to wen2023transformers for a detailed review of the large body of work and architectural variants of transformers that have been developed for time series applications.
@@
Attention in Vision The quadratic scaling in the input of the attention mechanism has made transformers a computationally prohibitive architecture for applications involving long sequences. Nonetheless, through techniques such as patching, the attention mechanism and transformers have also enjoyed success in computer vision, where high-resolution image data presents an increased computational cost. In fact, in vision transformers (ViT), introduced in dosovitskyi2021image, the input image is subdivided into patches; attention is then applied across the shorter sequence of patches, thus decreasing the overall cost.
@@
Neural Operators Given the success of the methodology in computer vision, there has been growing interest in extending the transformer methodology to the operator learning context. This operator setting concerns learning mappings between infinite dimensional spaces of functions. In this case, the domains on which the functions are defined (spatial and temporal) may inherently be a continuum. Neural operators (kovachki2023neural), generalizations of neural networks that learn operators mapping between infinite-dimensional spaces, have been developed for this task. As they parametrize operators acting on spaces of functions, the learned parameters of neural operators may be applied to any discretization of the input function; this property of neural operators is referred to as “discretization invariance”. Discretization invariance is desirable both at a theoretical and practical level. Indeed, for a neural network to approximate an operator acting on functions defined on a continuum, its parameters cannot depend on the discretizations of the input functions that are used for learning the parameters themselves. On the other hand, discretization invariance allows to scalably train models on coarser discretizations of functions and to deploy or finetune trained models at different resolutions for downstream tasks. A recent example of the success of this approach is the GenCast medium-range weather forecasting model (Price2025), which because of discretization invariance properties is pretrained on low resolutions and finetuned at the desired predictive resolution.
@@
Neural Operators. An integral kernel interpretation of the attention mechanism appearing for example in tsai2019transformer; martins2020sparse; moreno2022kernel; wright2021transformers; cao2021choose; kovachki2023neural inspired a range of attention-based architectures for operator learning. For example, the work of cao2021choose leverages insight from the integral kernel formulation to design novel attention mechanisms that do not employ the softmax function. The author designs attention mechanisms using integral kernels which may be approximated by a Fourier-type projection, leading to quadratic complexity, and by a Galerkin projection, leading to linear complexity. While the resulting attention maps are discretization invariant, the full architecture proposed uses convolutional neural networks to downsample the input to a feature map of sufficiently small resolution for the method to be scalable. Hence the full architecture is not a neural operator. Furthermore, the lack of the softmax nonlinearity leads to a mechanism lacking the probabilistic description outlined in this paper, thus differing from the original definition of attention in vaswani2017attention. In li2024transformer the authors take the Fourier attention approach as in cao2021choose to design a neural operator for large eddy simulation, while in luo2024hierarchical the authors take the Galerkin attention approach from cao2021choose to design a novel neural operator architecture. In li2023transformer the authors present “OFormer” a neural operator with an encoder-decoder structure that employs cross-attention to obtain the latent embeddings and then stacked self-attention blocks; the model uses the attention from cao2021choose and random Fourier features for query coordinate encodings while using a recurrent multi-layer perceptron for unrolling in time-dependent problems as opposed to masked attention in vaswani2017attention. In the recent work rahman2024pretraining, the authors propose a neural operator architecture that employs an attention mechanism that acts across channels of the function in the latent embedding, and do so in the context of pre-training a foundation model for PDE solution operators. We demonstrate how such an architecture can be derived from our formulation in Remark 16. As in our paper, various attention-related concepts are defined in a function-space setting, but our formulations are more general; this issue is discussed in detail at the relevant point in the paper. We note that the architectures introduced in this paper rely solely on the attention mechanism defined in the continuum, hence they are neural operators; furthermore, because of the simple design, with a slight modification to our architectures we may prove the first universal approximation theorem for transformer operators.
@@
Non discretization invariant architectures. In guibas2022efficient the authors propose a modification of the Fourier neural operator that involves mixing of tokens in Fourier space. This is achieved by first applying the FFT to the sequence of tokens and then applying a MLP to each token, before inverting back to the original space. This approach differs fundamentally from the original definition of attention in vaswani2017attention and lacks the probabilistic description outlined in this paper. Furthermore, because of learnable positional embeddings and patching via convolution, the AFNO is not discretization invariant. Various other approaches have been developed. The work hao2023gnot introduces “GNOT”, an architecture employing a normalized attention layer that handles multiple inputs. This design, along with a learned gating mechanism for input coordinates, allows the model to be applied to irregular meshes and multi-scale problems; however, this architecture employs an attention mechanism that does not involve rescaling by the grid, hence not directly allowing for discretization invariance. In ovadia2023vito the authors propose to use an architecture composed of a ViT in the latent space of a U-Net (ronneberger2015unet). The proposed model takes as input the input field evaluated on a grid with associated coordinate values and outputs an approximation of the solution to PDE inverse problems. While it is shown that the architecture achieves state-of-the-art accuracy, the parameters in the scheme are not invariant to the input function resolution. In wang2025continuous the authors design “Continuous ViT”, an architecture consisting of a ViT encoder and a cross-attention decoder. The method handles input functions defined on space and time and the method is shown to present evidence of discretization invariance, with test discretizations close to training discretizations. However, while the architecture is discretization invariant in time, it is not by design in space. This is because the tokenization procedure employed by standard ViTs is not discretization invariant; we elaborate on this point in Remark 15.
@@
Transformers for Continuum Data The use of attention-based transformer models in the context of dynamical systems and data assimilation problems, where spatial and temporal domains of the systems are continuous, is at its infancy. For a comprehensive review on the application of machine learning to data assimilation we refer the reader to cheng2023machine. In this context, the efficacy of transformers has been explored for forecasting in numerical weather prediction. In chattopadhyay2022towards, a transformer module is integrated in the latent space of a U-Net (ronneberger2015unet) architecture in order to leverage the permutation equivariance property of the transformer to improve physical consistency and forecast accuracy. The work in bi2023accurate achieves state-of-the-art performance among end-to-end learning methods for numerical weather prediction by applying a ViT-like architecture and employing the shifted window methodology from liu2021swin. These approaches, however, are not discretization invariant by design.
@@
Mathematics of Transformers Mathematical foundations for the methodology are only now starting to emerge, however most analysis has been preformed in the finite-dimensional setting. In the recent work geshkovski2023emergence, the authors study the attention dynamics by viewing tokens as particles in an interacting particle system. The formulation of the continuous-time limit of the dynamics, where the limit is taken in “layer time” allows to investigate the emergence of clusters among tokens. In Yun2020Are, the authors prove that transformers are universal approximators of continuous permutation equivariant sequence-to-sequence functions with compact support. More broadly, developing a firm theoretical framework for the subject is necessary to understand the empirical performance of the methodology and to inform the design of novel attention-based architectures.
@@ 2 Continuum Attention
In this section we formulate a continuum analogue of the attention mechanism described in vaswani2017attention. We note that in the transformer architecture of vaswani2017attention, a distinction is made between the self-attention and cross-attention functions. The former, a particular instance of (1) where $u=v$, can be used to produce a representation of the input that captures intra-sequence correlations. On the other hand, cross-attention is used to output a representation of the input $u$ that models cross-correlations between both inputs $u$ and $v$. Indeed, in vaswani2017attention cross-attention is employed in the decoder component of the transformer architecture so that the input to the decoder attends to the output of the encoder. For natural language applications in particular, the form of the cross-attention function has allowed for sequences of arbitrary length $N$ to be outputted from a transformer architecture while attending to sequences of different lengths, e.g. as outputs of length $M$ from a self-attention encoder block.
For pedagogical purposes we subdivide the presentation between the self-attention operator and the cross-attention operator, which are treated in Subsection 2.1 and Subsection 2.2 respectively. In Subsection 2.1.1 we define self-attention for discrete sequences indexed on finite bounded sets. Here, we outline the interpretation of self-attention as an expectation under a family of carefully defined discrete probability measures, each acting on the space of sequence indices. We show that under this formulation we recover the standard definition of self-attention from vaswani2017attention. We proceed in Subsection 2.1.2 by formulating self-attention as an operator mapping between spaces of functions defined on bounded uncountable sets. This interpretation is natural, as functions may be thought of as “sequences” indexed over an uncountable domain. As in the discrete case, we formulate the self-attention operator as an expectation under a carefully defined probability measure acting on the domain of the input. We obtain an approximation result given by Theorem 6 that shows that for continuous functions the self-attention mechanism as implemented in practice may be viewed as a Monte Carlo approximation of the self-attention operator described. We mimic these constructions in Subsections 2.2.1 and 2.2.2 for the discrete and continuous formulations of cross-attention, respectively, and obtain an analogous approximation result given by Theorem 12. As in (1), throughout this section $Q\in\mathbbm{R}^{d_{K}\times d_{u}},K\in\mathbbm{R}^{d_{K}\times d_{v}},V\in\mathbbm{R}^{d_{V}\times d_{v}}$ are learnable weights that parametrize the attention operators. When we consider self-attention we will have $d_{v}=d_{u}.$
@@ 3 Continuum Patched Attention
$\widetilde{u}(p)={\left.\kern-1.2ptu\mathchoice{\vphantom{\big|}}{}{}{}\right|_{D_{p}}},$ (6)
@@ 3.1 Patched Self-Attention
An image may be viewed as a mapping $u:D\to\mathbbm{R}^{d_{u}}$, where $D\subset\mathbbm{Z}^{2}$, and where $d_{u}=1$ for grey-scale and $d_{u}=3$ for RGB-valued images. In the context of vision transformers (dosovitskyi2021image), the input to the attention mechanism is a sequence of $P$ “flattened” patches of the input image; we next outline the details to this procedure.
Letting $N$ be the total number of discretization points in $D$ and $P$ the number of patches, in the discrete setting each function patch $\widetilde{u}(p):D^{\prime}\to\mathbbm{R}^{d_{u}}$ may be represented as an $\mathbbm{R}^{N/P\times d_{u}}$ matrix. The flattening procedure of each patch in ViT involves reshaping this matrix into an $\mathbbm{R}^{N/P\cdot d_{u}}$ vector. The full image is thus represented as a sequence of $P$ vectors of dimension $N/P\cdot d_{u}$. Each patch vector is then linearly lifted to an embedding space of dimension $d_{\textrm{model}}$. The embedded image, to which attention is applied, may hence be viewed as a matrix $\widetilde{u}\in\mathbbm{R}^{P\times d_{\textrm{model}}}$. Therefore, ViT may be cast in the setting of the discrete interpretation of the self-attention operator outlined in Subsection 2.1.1, where the domain on which attention is applied is $D^{P}$. A key observation to the ensuing discussion is that the application of a linear transformation to a “flattened” patch in ViT breaks the mesh-invariance of the architecture; however, this operation may be viewed as a nonlocal transformation on each patch. We leverage this insight in Section 4 to design discretization invariant transformer neural operators that employ patching. We further note that in practice, patching in ViT is often also achieved via strided convolutions, with stride equivalent to the kernel size; however, because the kernel size determines the number of parameters, such a procedure also breaks discretization invariance as it leads to architectures with parameters dependent on resolution of the input.
@@
The work of rahman2024pretraining employs an attention mechanism that acts across the channels of the function $u\in\mathcal{U}\bigl(D;\mathbbm{R}^{d_{\textrm{model}}}\bigr)$ in the latent space of a neural operator architecture. Indeed the “codomain” attention mechanism defined may be viewed as a particular variant of Definition 14, where $D^{\prime}\cong D$, where $\mathsf{A}_{\mathsf{P}}(\widetilde{u})(j)$ is defined for $j\in\llbracket 1,d_{\textrm{model}}\rrbracket$ and where $\mathsf{Q},\mathsf{K},\mathsf{V}$ are implemented as integral operators. In particular the indexing is by channels and not by patch. For problems in computer vision, similar ideas have been explored with discrete versions of the $\mathsf{Q},\mathsf{K},\mathsf{V}$ operators (chen2017sca).